# algorithmic

### v2

@@ 2. Unknown
[…]
@@
$\cdot\mathbb{E}_{\pi^{*}}\left[\sum^{H}_{h=1}{(N_{h}(s_{h},a_{h})+1)}^{-1/2}\right]\Big{)}$
@@
Kong & Yang (2022) study the problem of optimal policy learning from critique feedback (see Section 3.2), i.e., binary rewards for state-action pairs, with as few queries to the human as possible. They assume an underlying ground-truth human feedback model that leads to a positive evaluation for a state-action pair if it exceeds a specific threshold evaluated at that pair. In addition, the learning process consists of two phases: First, exploring the environment in an unsupervised manner, and then querying user feedback in an active reward learning phase to learn the human feedback model. This learning process is again analyzed in two variants: Either the exploration phase was performed externally, and a data set consisting of trajectories is provided (offline), or this data set is actively collected itself (online). For both variants, an active learning algorithm is proposed that essentially selects query points (state-action pairs) that provide the most information gain given the points already designated to be queried. For the online variant, an exploration strategy based on optimistic least-squares value iteration (Jin et al., 2020) is also introduced for tabular or linear MDPs. In both variants, policy learning is carried out by a pessimistic value iteration with the empirical transitions and the estimated reward function, resulting in UCBVI-Planning (online) and LCBVI-Tabular-Offline (offline). Under the assumption of bounded noise (Massart & Nédélec, 2006) or low-noise assumption (Korba et al., 2017; Haddenhorst et al., 2021), bounds on the performance gap of both algorithms are derived.
@@
There have been some theoretical analyses regarding the question in how far, or if at all, preference-based feedback in the form of trajectory comparisons is more suitable compared to numerical feedback. Ji et al. (2023c) suggest a human rating model for this purpose in the numerical feedback case and analyze the LCB algorithm (Jin et al., 2021) in order to compare it with the pessimistic MLE (Zhu et al., 2023a). It is shown that under specific assumptions, LCB has a constant performance gap, while the preference-based pessimistic MLE under similar assumptions has a similar bound as in Table 7.
Wang et al. (2023b) provide reduction-based algorithms that can directly utilize state-of-the-art results in reward-based RL for RLHF with utility-based and general state-action and trajectory-based comparison feedback. They show, in general, how theoretical results of the underlying standard RL algorithm can be translated to theoretical results for the resulting preference-based RL algorithm. For some special cases, such as MDPs with finite Eluder dimension and utility-based preference feedback, the theoretical guarantees are explicitly derived using state-of-the-art RL algorithms that are qualitatively similar to explicit preference-based RL algorithms.
@@ 7.3 Nash Learning from Human Feedback
The majority of theoretical works use the modeling of (pairwise comparison) feedback by means of a link function (see Section 5.1.4). Even if this often leads to simpler derivations, this modeling has the decisive disadvantage that it imposes a transitivity of the human feedback that does not necessarily prevail in reality. In other words, it is quite possible that preference cycles can occur. For this reason, there is a new direction in theoretical work that dispenses with parametric modeling of the preference probability similar to Chen et al. (2022) but uses it to formulate a new learning objective. Specifically, the problem is considered from a game theory perspective, where two policies each propose a trajectory that should be highly preferred by the human user. Thus, the goal is to find a policy that suggests trajectories that are preferred to the trajectories of any other policy, i.e., a Nash equilibrium or a von Neumann winner.
This learning variant was first considered by Wang et al. (2023b), who showed that the problem can be reduced to finding restricted Nash equilibria in a multi-agent RL problem (based on numerical rewards). For special situations of the latter problem, wrapper algorithms are proposed that have been shown to find the von Neumann winner with high probability. The learning problem was recently taken up and analyzed by Munos et al. (2023) and Ye et al. (2024) in a KL-regularization variant. While the former considers the online learning setting assuming a known preference model, the latter considers both online as well as offline learning settings and the preference model belonging to a finite function class.
@@ 8 Applications and Benchmarks
The field of RLHF has advanced significantly in the last few years, with increasing interest driven by prominent applications. First and foremost are applications to large language models, exemplified by ChatGPT (OpenAI, 2022). This section starts by providing a sample of such applications, showcasing how this technology is being utilized in fields as varied as robotics, language processing, image generation, and more. We will also delve into libraries that provide foundational support for RLHF research, enabling researchers and practitioners to experiment with and refine a range of approaches. We then explore a spectrum of benchmarks that have been developed to standardize and simplify the evaluation of new approaches, offering insights into their performance in different settings. Finally, and closely related to those benchmarks, we will discuss common evaluation practices.
@@ 8.1 Applications
There is a long history of using control environments as benchmark tasks for RL. In addition to the breadth of available environments, control applications are of particular interest because tasks are often hard to specify. Christiano et al. (2017) demonstrated the effectiveness of RLHF in games as well as simulated continuous control tasks, matching the performance of RL agents trained on ground-truth rewards with a fraction of the feedback. Extending to robotics, Ding et al. (2023) trained a reward model for diverse tasks with a single robot, achieving human-like behavior. Kupcsik et al. (2018) applied RLHF for precise robot-to-human handovers. Similarly, Abramson et al. (2022) used RLHF in the Playhouse simulator, a platform for sensorimotor task training, and Milani et al. (2022) showcase an application in the context of the MineRL Basalt competition for Minecraft tasks. Recently, Dong et al. (2024) use RLHF to guide a diffusion-based planning model.
@@
Generative models, i.e., models that generate new data instead of just predicting labels, can be framed as an RL setting in which a policy assembles the output through its actions. In the context of language models, this means that the language model is interpreted as a policy with tokens as actions. Using this reframing, we can use RLHF approaches to fine-tune generative models to produce preferred outputs. ChatGPT (OpenAI, 2022) and GPT-4 (OpenAI, 2023) are prime examples of language models fine-tuned using RLHF. These applications build on earlier work, such as by Ouyang et al. (2022), Ziegler et al. (2020) and Glaese et al. (2022). This method extends to text summarization (Gao et al., 2018; 2020; Stiennon et al., 2020), dialogue summarization (Chen et al., 2023), and question answering (Nakano et al., 2022). In image generation, Lee et al. (2023) and Xu et al. (2023) demonstrate the use of reward modeling for text-to-image tasks, while Pinto et al. (2023) and Kazemi et al. (2020) explore RLHF applications in broader computer vision tasks. Interestingly, in the context of LLMs, reward learning has also been expressed as density estimation (Dumoulin et al., 2024) instead of the supervised approach described in Section 5.
@@
In the context of recommender systems, Xue et al. (2023b) have shown the potential of RLHF in optimizing for long-term engagement. Although it is, in principle, possible to algorithmically evaluate policies in this domain, these rewards are sparse. To combat this, Xue et al. (2023b) use RLHF to distill sparse, global feedback into a dense reward model.
These diverse applications underscore RLHF’s adaptability and its growing importance in various technological domains, paving the way for innovative solutions and enhanced human-computer interactions.
@@ 8.2 Supporting Libraries
Several libraries have emerged that aim to provide a toolset for implementing and experimenting with RLHF and reward learning algorithms, contributing to the ease and efficiency of research and development. One notable example is the imitation library (Gleave et al., 2022b). It encompasses a collection of imitation and reward learning algorithms, including those introduced in the seminal work by Christiano et al. (2017). In the offline realm, Clean-Offline-RLHF (Yuan et al., 2024) provides implementations for offline RL algorithms with human feedback. Two other libraries, APReL (Bıyık et al., 2022b) and POLAR (Tucker et al., 2022), focus on the Bayesian setting. Bıyık et al. (2022b) provide a specialized framework for preference-based reward learning with a focus on Bayesian methods. Meanwhile, Tucker et al. (2022) introduce a framework designed for Bayesian reward learning from multiple feedback types, including pairwise preferences, in MATLAB. Finally, the domain of language model fine-tuning, the trlX library (Castricato et al., 2023) offers a toolkit specifically designed for language model training. It specializes in the fine-tuning of transformer-based language models, treating the language model as the policy in an RLHF setup.
@@ 8.3 Benchmarks
In the domain of language modeling, Truthful QA (Lin et al., 2022) serves as a benchmark that measures the truthfulness of models. Also, in the context of language models, Ramamurthy et al. (2023) introduce a set of pre-trained reward models, learned from human feedback, as benchmarks. These models serve as reference points for evaluating new RLHF techniques against established standards.
@@ 8.4 Datasets
Since language model fine-tuning is a popular application of RLHF and generally does not iteratively refine the reward model, many datasets have been developed for this purpose. Particularly notable are hh-rlhf (Bai et al., 2022a) and PKU-Safe-RLHF (Ji et al., 2023a), two datasets focusing on harmless and helpful responses, the OpenAssistant datasets (oasst1, oasst2(Köpf et al., 2023), containing not only response rankings but also ratings on various dimensions, the summarize_from_feedback dataset (Stiennon et al., 2020) focusing on preferences over text summaries, the Stanford Human Preferences Dataset (SHP(Ethayarajh et al., 2022), which is based on Reddit responses, the WebGPT dataset (webgpt_comparisons(Nakano et al., 2022), focused on long-form question answering and the HelpSteer (Wang et al., 2023c) dataset, which is not based on preferences but instead gives ratings on for 4 attributes (helpfulness, correctness, coherence, complexity) for each response.
@@ 8.5 Evaluation
Assessing learned behavior is crucial for the evaluation of an RLHF system. In domains with ground-truth rewards, these can be used for policy evaluation (Christiano et al., 2017). However, many RLHF applications lack this clarity. Ouyang et al. (2022), for instance, evaluate the quality of language model responses by having labelers rate the output quality on a test set of prompts, highlighting the significance of human judgment in assessing model outputs. Jain et al. (2015) use direct Likert-scale scores for evaluations, including self-assessments by trainers and cross-evaluations by others. Losey et al. (2022) extend this with a Likert-scale survey and free-form participant comments, comparing evaluations based on known true rewards with subjective experiences. Moreover, Abramson et al. (2022) employ a multi-stage evaluation scheme that includes scripted probe tasks, a standardized test suite evaluated by humans, and full interactive assessments, demonstrating the need for diverse and thorough evaluation methodologies in RLHF.
@@
Direct reward model evaluation complements policy assessment. While reward model accuracy is a more direct measure of preference-learning success, the ultimate goal is inducing effective policies. A perfectly accurate reward model is often not necessary to induce a good policy, which is the actual goal of RLHF. Therefore, both evaluation methods are ideally used in combination. Jain et al. (2015) also use a ranking loss method for test sets of trajectories, compared against expert evaluations with known Likert-scores. This approach provides quantitative measures of the reward model’s fidelity. In addition, Wilde & Alonso-Mora (2022) compare parameter-based and reward-based evaluation measures for learned reward functions, identifying strengths and weaknesses in both methods and contributing to a more nuanced understanding of reward model assessment in RLHF. These approaches provide a quantitative measure of the reward model’s accuracy in reflecting human preferences and expert judgments. For a detailed discussion of reward model evaluation, also refer to Section 5.3.
@@ 9 Discussion and Conclusion
In this survey, we have provided an overview of the current state of RLHF, highlighting its evolution from PbRL and examining its broad applications across various domains like control, natural language processing, and computer vision. While our survey captures the current state and many significant trends and advancements in RLHF, we acknowledge the rapid expansion of this field and the inevitable limitations in covering every extension and application in depth. We will discuss some of these extensions, open questions, and conclusions in this section.
We have specifically focused on RLHF methods where a reward function is learned online from human feedback. There have been some recent works that are outside of this scope and yet propose promising new methods to learn human-aligned objectives, such as offline learning of reward functions (Shin et al., 2023) or privacy-preserving alignment based on differential privacy (Wu et al., 2024).
@@
An alternative approach to RLHF is to learn objectives from a pre-trained AI system instead of human feedback. This has been termed RL from AI feedback (RLAIF) and leverages foundation models as a source of preference and has shown successful for language-model fine-tuning (Bai et al., 2022b; Sun et al., 2024), generating intrinsic motivation for text-based games (Klissarov et al., 2024), as well as learning rewards (Wang et al., 2024b) or coding reward functions (Ma et al., 2024; Xie et al., 2024) for control tasks without any human involvement. Closely related is the setting of assisted evaluation, studied by Saunders et al. (2022), where a language model is used to generate critiques of language model outputs, thereby assisting human evaluators.
@@
Most work on RLHF implicitly assumes that tasks can be specified by maximization of expected accumulated scalar rewards. This assumption, called the reward hypothesis (Silver et al., 2021), is under active debate (Lambert, 2021; Vamplew et al., 2022; Bowling et al., 2023; Skalse & Abate, 2022) in the RL community. On a related note, Skalse & Abate (2024) investigate the sensitivity of inverse RL to this misspecification. Recent approaches in RLHF are, for instance, considering more complex objective functions, such as multi-objective frameworks involving non-linear aggregation of expected accumulated vector rewards (Qian et al., 2023).
@@
Many more extensions of RLHF are inspired by revisiting classic RL topics under the RLHF lens. This is exemplified by studies on exploration (Liang et al., 2022), reward feature learning (Katz et al., 2021), reward shaping (Xiao et al., 2020), multi-task RL (Ouyang et al., 2022; Abramson et al., 2022; Myers et al., 2023), hierarchical RL (Pinsler et al., 2018), hindsight experience replay (Zhang et al., 2023), risk-sensitive RL (Chen et al., 2024), safe RL (Dai et al., 2024; Cosner et al., 2022), fair RL (Siddique et al., 2023), or continual learning (Zhang et al., 2024). As discussed in Section 4.2, the intersection of RLHF with HCI also offers a fertile ground for future research, especially for refining feedback mechanisms. It is crucial to keep human psychology in mind when designing these systems and to learn from other related fields that already studied such issues extensively.
In addition to those extensions, current RLHF methods also have challenges and limitations to be aware of. Without assistance during feedback, it is limited by the tasks humans can reliably judge (Leike, 2022; Leike et al., 2018; Wu et al., 2021; Christiano et al., 2018). This challenge can be seen when trying to fine-tune a LLM to give factual answers, where humans often prefer assertive, but wrong responses (Hosking et al., 2024). As another limitation, current RLHF approaches often fail to learn the actual causes of human feedback, learning correlations instead (Tien et al., 2022). The common separation between the policy and the reward model also limits how the agent can reason about its knowledge of the human preferences, a limitation that may be lifted by tighter integration such as in the cooperative inverse RL setting (also called assistance games) (Hadfield-Menell et al., 2016; Shah et al., 2021a). Additionally, the RLHF framework is limited in how it can reason about its knowledge of human preferences Casper et al. (2023) offer a thorough analysis of further issues and limitations of RLHF, highlighting the practical constraints of current approaches. Adding to that, from a theoretical perspective, a primary challenge lies in further relaxing underlying assumptions. This requires striking a delicate balance: On the one hand, ensuring the assumptions are not overly restrictive to encompass a broad range of practical use cases, and on the other, maintaining the feasibility of theoretical guarantees for computationally efficient algorithms. Key questions in this context are whether it is possible to design algorithms that do not need to actively maintain a policy space and eliminate sub-optimal policies nor rely on a computation oracle. Recent work such as Wang et al. (2023b) or Wu & Sun (2024) give hope that this may be possible.

### v3

@@ 2. Unknown
$\mathcal{O}\left(\tilde{d}He^{H}|\mathcal{A}|\sqrt{\mu\log(nR_{\mathrm{max}}H/\delta)}\right)$ $\mu\text{-finite spectrum, }$
$\mathcal{O}\left(\tilde{d}He^{H}|\mathcal{A}|\sqrt{{(\log(nR_{\mathrm{max}}H)/\delta)}^{1+1/\mu}}\right)$ $\mu\text{-exponential decay, }$
$\mathcal{O}\left(\tilde{d}He^{H}|\mathcal{A}|{(nR_{\mathrm{max}})}^{\kappa^{*}}\sqrt{\log(nR_{\mathrm{max}}H/\delta)}\right)$ $\mu\text{-polynomial decay,}$
@@
$\hphantom{\Big(}\cdot\mathbb{E}_{\pi^{*}}\left[\sum^{H}_{h=1}{(N_{h}(s_{h},a_{h})+1)}^{-1/2}\right]\Big)$
@@
Kong & Yang (2022) study the problem of optimal policy learning from critique feedback (see Section˜3.2), i.e., binary rewards for state-action pairs, with as few queries to the human as possible. They assume an underlying ground-truth human feedback model that leads to a positive evaluation for a state-action pair if it exceeds a specific threshold evaluated at that pair. In addition, the learning process consists of two phases: First, exploring the environment in an unsupervised manner, and then querying user feedback in an active reward learning phase to learn the human feedback model. This learning process is again analyzed in two variants: Either the exploration phase was performed externally, and a data set consisting of trajectories is provided (offline), or this data set is actively collected itself (online). For both variants, an active learning algorithm is proposed that essentially selects query points (state-action pairs) that provide the most information gain given the points already designated to be queried. For the online variant, an exploration strategy based on optimistic least-squares value iteration (Jin et al., 2020) is also introduced for tabular or linear MDPs. In both variants, policy learning is carried out by a pessimistic value iteration with the empirical transitions and the estimated reward function, resulting in UCBVI-Planning (online) and LCBVI-Tabular-Offline (offline). Under the assumption of bounded noise (Massart & Nédélec, 2006) or low-noise assumption (Korba et al., 2017; Haddenhorst et al., 2021), bounds on the performance gap of both algorithms are derived.
@@
There have been some theoretical analyses regarding the question of how far, or if at all, preference-based feedback in the form of trajectory comparisons is more suitable compared to numerical feedback. Ji et al. (2023c) suggest a human rating model for this purpose in the numerical feedback case and analyze the LCB algorithm (Jin et al., 2021) in order to compare it with the pessimistic MLE (Zhu et al., 2023a). It is shown that under specific assumptions, LCB has a constant performance gap, while the preference-based pessimistic MLE under similar assumptions has a similar bound as in Table˜5.
Wang et al. (2023b) provide reduction-based algorithms that can directly use state-of-the-art results in reward-based RL for RLHF with utility-based and general state-action and trajectory-based comparison feedback. They show, in general, how theoretical results of the underlying standard RL algorithm can be translated to theoretical results for the resulting preference-based RL algorithm. For some special cases, such as MDPs with finite Eluder dimension and utility-based preference feedback, the theoretical guarantees are explicitly derived using state-of-the-art RL algorithms that are qualitatively similar to explicit preference-based RL algorithms.
@@ 7.3 Nash Learning from Human Feedback
The majority of theoretical works use the modeling of (pairwise comparison) feedback by means of a link function (see Section˜5.1.4). Even if this often leads to simpler derivations, this modeling has the decisive disadvantage that it imposes transitivity of the human feedback that does not necessarily prevail in reality. In other words, it is quite possible that preference cycles can occur. For this reason, there is a new direction in theoretical work that dispenses with parametric modeling of the preference probability similar to Chen et al. (2022) but uses it to formulate a new learning objective. Specifically, the problem is considered from a game theory perspective, where two policies each propose a trajectory that should be highly preferred by the human user. Thus, the goal is to find a policy that suggests trajectories that are preferred to the trajectories of any other policy, i.e., a Nash equilibrium or a von Neumann winner.
This learning variant was first considered by Wang et al. (2023b), who showed that the problem can be reduced to finding restricted Nash equilibria in a multi-agent RL problem (based on numerical rewards). For special situations of the latter problem, wrapper algorithms are proposed that have been shown to find the von Neumann winner with high probability. The learning problem was recently taken up and analyzed by Munos et al. (2024) and Ye et al. (2024) in a KL-regularization variant. While the former considers the online-learning setting assuming a known preference model, the latter considers both online as well as offline learning settings and the preference model belonging to a finite function class.
@@ 8 Applications and Benchmarks
The field of RLHF has advanced significantly in the last few years, with increasing interest driven by prominent applications. First and foremost are applications to large language models, exemplified by ChatGPT (OpenAI, 2022). This section starts by providing a sample of such applications, showcasing how this technology is being used in fields as varied as robotics, language processing, image generation, and more. We will also delve into libraries that provide foundational support for RLHF research, enabling researchers and practitioners to experiment with and refine a range of approaches. We then explore a spectrum of benchmarks that have been developed to standardize and simplify the evaluation of new approaches, offering insights into their performance in different settings. Finally, and closely related to those benchmarks, we will discuss common evaluation practices.
@@ 8.1 Applications
There is a long history of using control environments as benchmark tasks for RL. In addition to the breadth of available environments, control applications are of particular interest because tasks are often hard to specify. Christiano et al. (2017) demonstrated the effectiveness of RLHF in games as well as simulated continuous control tasks, matching the performance of RL agents trained on ground-truth rewards with a fraction of the feedback. Extending to robotics, Ding et al. (2023) trained a reward model for diverse tasks with a single robot, achieving human-like behavior. Kupcsik et al. (2018) applied RLHF for precise robot-to-human handovers. Similarly, Abramson et al. (2022) used RLHF in the Playhouse simulator, a platform for sensorimotor task training, and Milani et al. (2022) showcase an application in the context of the MineRL Basalt competition for Minecraft tasks. Recently, Dong et al. (2024b) use RLHF to guide a diffusion-based planning model.
@@
Generative models, i.e., models that generate new data instead of just predicting labels, can be framed as an RL setting in which a policy assembles the output through its actions. In the context of language models, this means that the language model is interpreted as a policy with tokens as actions. Using this reframing, we can use RLHF approaches to fine-tune generative models to produce preferred outputs. ChatGPT (OpenAI, 2022) and GPT-4 (OpenAI, 2023) are prime examples of language models fine-tuned using RLHF. These applications build on earlier work, such as by Ouyang et al. (2022), Ziegler et al. (2020) and Glaese et al. (2022). This method extends to text summarization (Gao et al., 2018; 2020; Stiennon et al., 2020), dialogue summarization (Chen et al., 2023), and question answering (Nakano et al., 2022; Menick et al., 2022). In image generation, Lee et al. (2023) and Xu et al. (2023) demonstrate the use of reward modeling for text-to-image tasks, while Pinto et al. (2023) and Kazemi et al. (2020) explore RLHF applications in broader computer vision tasks. Interestingly, in the context of LLMs, reward learning has also been expressed as density estimation (Dumoulin et al., 2024) instead of the supervised approach described in Section˜5.
@@
In the context of recommender systems, Xue et al. (2023) have shown the potential of RLHF in optimizing for long-term engagement. Although it is, in principle, possible to algorithmically evaluate policies in this domain, these rewards are sparse. To combat this, Xue et al. (2023) use RLHF to distill sparse, global feedback into a dense reward model.
These diverse applications underscore the adaptability of RLHF and its growing importance in various technological domains, paving the way for innovative solutions and enhanced human-computer interactions.
@@ 8.2 Supporting Libraries
Several libraries have emerged that aim to provide a toolset for implementing and experimenting with RLHF and reward learning algorithms, contributing to the ease and efficiency of research and development. One notable example is the imitation library (Gleave et al., 2022b). It encompasses a collection of imitation and reward learning algorithms, including those introduced in the seminal work by Christiano et al. (2017). In the offline realm, Clean-Offline-RLHF (Yuan et al., 2024) provides implementations for offline RL algorithms with human feedback. Two other libraries, APReL (Bıyık et al., 2022b) and POLAR (Tucker et al., 2022), focus on the Bayesian setting. Bıyık et al. (2022b) provide a specialized framework for preference-based reward learning with a focus on Bayesian methods. Meanwhile, Tucker et al. (2022) introduce a framework designed for Bayesian reward learning from multiple feedback types, including pairwise preferences, in MATLAB. Finally, in the domain of language model fine-tuning, the trlX library (Castricato et al., 2023) offers a toolkit specifically designed for language model training. It specializes in the fine-tuning of transformer-based language models, treating the language model as the policy in an RLHF setup.
@@ 8.3 Benchmarks
In the domain of language modeling, Truthful QA (Lin et al., 2022) serves as a benchmark that measures the truthfulness of models. Also, in the context of language models, Ramamurthy et al. (2023) introduce a set of pretrained reward models, learned from human feedback, as benchmarks. These models serve as reference points for evaluating new RLHF techniques against established standards.
For evaluating language reward models specifically, Lambert et al. (2024) introduce RewardBench, a benchmark designed to test reward model performance across chat, safety, reasoning, and preference learning tasks. The benchmark includes challenging test cases such as adversarial prompts and edge cases where reward models commonly fail, providing standardized evaluation for both classifier-based and DPO-trained reward models. RewardBench has become a widely-used evaluation framework with a public leaderboard that tracks performance of submitted reward models.
@@ 8.4 Datasets
Since language model fine-tuning is a popular application of RLHF and generally does not iteratively refine the reward model, many datasets have been developed for this purpose. Particularly notable are hh-rlhf (Bai et al., 2022a) and PKU-Safe-RLHF (Ji et al., 2023a), two datasets focusing on harmless and helpful responses, the OpenAssistant datasets (oasst1, oasst2(Köpf et al., 2023), containing not only response rankings but also ratings on various dimensions, the summarize_from_feedback dataset (Stiennon et al., 2020) focusing on preferences over text summaries, the Stanford Human Preferences Dataset (SHP(Ethayarajh et al., 2022), [^10] This paper introduces the V-usable information framework for measuring dataset difficulty. The SHP dataset was created after the paper’s publication, using its techniques to identify learnable preference pairs from Reddit data. which is based on Reddit responses, the WebGPT dataset (webgpt_comparisons(Nakano et al., 2022), focused on long-form question answering and the HelpSteer (Wang et al., 2024d) dataset, which is not based on preferences but instead gives ratings on for 4 attributes (helpfulness, correctness, coherence, complexity) for each response.
@@ 8.5 Evaluation
Assessing learned behavior is crucial for the evaluation of an RLHF system. In domains with ground-truth rewards, these can be used for policy evaluation (Christiano et al., 2017). However, many RLHF applications lack this clarity. Ouyang et al. (2022), for instance, evaluate the quality of language model responses by having labelers rate the output quality on a test set of prompts, highlighting the significance of human judgement in assessing model outputs. Jain et al. (2015) use direct scores on a Likert scale for evaluations, including self-assessments by trainers and cross-evaluations by others. Losey et al. (2022) extend this with a survey based on scores and free-form participant comments, comparing evaluations based on known true rewards with subjective experiences. Moreover, Abramson et al. (2022) employ a multi-stage evaluation scheme that includes scripted probe tasks, a standardized test suite evaluated by humans, and full interactive assessments, demonstrating the need for diverse and thorough evaluation methodologies in RLHF.
@@
Direct reward model evaluation complements policy assessment. While reward model accuracy is a more direct measure of preference-learning success, the ultimate goal is inducing effective policies. A perfectly accurate reward model is often not necessary to induce a good policy, which is the actual goal of RLHF. Therefore, both evaluation methods are ideally used in combination.
Jain et al. (2015) also use a ranking loss method for test sets of trajectories, compared against expert evaluations with known scores. This approach provides quantitative measures of the reward model’s fidelity. In addition, Wilde & Alonso-Mora (2022) compare parameter-based and reward-based evaluation measures for learned reward functions, identifying strengths and weaknesses in both methods and contributing to a more nuanced understanding of reward model assessment in RLHF. These approaches provide a quantitative measure of the reward model’s accuracy in reflecting human preferences and expert judgements. For a detailed discussion of reward model evaluation, also refer to Section˜5.3.
@@ 9 Discussion and Conclusion
In this survey, we have provided an overview of the current state of RLHF, tracing its evolution from PbRL and examining its broad applications across various domains like control, natural language processing, and computer vision. Given the rapid expansion of this field, our coverage necessarily has limitations in addressing every extension and application in full depth. In particular, given the rapidly evolving nature of LLM research, our coverage of techniques specific to this domain is less exhaustive than our treatment of control and robotics applications. In this concluding section, we discuss key extensions and alternative approaches, identify open questions, and highlight future research directions.
9.1 Extensions and Related Approaches
While our survey focuses on RLHF methods that learn reward functions online from human feedback, several promising approaches extend beyond this scope while addressing the same core challenge of learning human-aligned objectives.
@@ 9.1 Extensions and Related Approaches
An increasingly important alternative is RL from AI feedback (RLAIF), which replaces human feedback with evaluations from pre-trained AI systems. This approach leverages foundation models as preference sources and has demonstrated success across diverse applications: language model fine-tuning (Bai et al., 2022b; Sun et al., 2024), generating intrinsic motivation for text-based games (Klissarov et al., 2024), and learning rewards (Wang et al., 2024c) or coding reward functions (Ma et al., 2024b; Xie et al., 2024) for control tasks. A hybrid approach (assisted evaluation) involves AI systems assisting rather than replacing human evaluators, such as generating critiques of model outputs (Saunders et al., 2022) or enhancing responses with metadata such as citations (Nakano et al., 2022; Menick et al., 2022).
@@
While most work on RLHF implicitly assumes that tasks can be specified by maximization of expected accumulated scalar rewards, this reward hypothesis (Silver et al., 2021) is under active debate in the RL community (Lambert, 2021; Vamplew et al., 2022; Bowling et al., 2023; Skalse & Abate, 2022). Skalse & Abate (2024) investigate the sensitivity of inverse RL to reward misspecification, and recent RLHF approaches are beginning to move beyond scalar rewards through more complex objective functions, such as multi-objective frameworks with non-linear aggregation of vector rewards (Qian et al., 2023).
@@
Finally, the RLHF framework has inspired numerous extensions across RL domains that revisit classic RL topics through the lens of human feedback. These extensions span three main categories: fundamental algorithmic concerns such as exploration (Liang et al., 2022), reward feature learning (Katz et al., 2021), hindsight experience replay (Zhang et al., 2023a), and reward shaping (Xiao et al., 2020); advanced learning paradigms including multi-task learning (Ouyang et al., 2022; Abramson et al., 2022; Myers et al., 2023), continual learning (Zhang et al., 2024), and hierarchical RL (Pinsler et al., 2018); and specialized domains such as risk-sensitive (Chen et al., 2024), safe (Dai et al., 2024; Cosner et al., 2022), and fair RL (Siddique et al., 2023). Additionally, as discussed in Section˜4.2, the intersection of RLHF with human-computer interaction (HCI) offers rich opportunities for advancing feedback collection mechanisms. It is crucial to keep human psychology in mind when designing these systems and to learn from other related fields that already studied such issues extensively.
9.2 Current Limitations
Despite its successes, RLHF faces several limitations that constrain its broader applicability. Casper et al. (2023) offer a thorough analysis of open problems and fundamental limitations of RLHF; beyond that, we highlight several key limitations in the following.
A fundamental limitation concerns feedback quality and interpretation. Without assistance during feedback, RLHF is limited by the tasks humans can reliably judge (Leike, 2022; Leike et al., 2018; Wu et al., 2021; Christiano et al., 2018) – a problem exemplified in language model fine-tuning, where humans often prefer assertive but incorrect responses (Hosking et al., 2024). Furthermore, current approaches often fail to learn the actual causes of human feedback, learning spurious correlations instead (Tien et al., 2022).
Current approaches also struggle with awareness of gaps in preference understanding. The common separation between the policy and the reward model limits how the agent can reason about its knowledge of human preferences. This limitation may be addressed by tighter integration such as in the cooperative inverse RL setting (also called assistance games) (Hadfield-Menell et al., 2016; Shah et al., 2021a), where the agent simultaneously learns about human preferences and acts to maximize them, enabling behaviors like conditional planning while awaiting feedback and relevance-aware active learning. More broadly, the limited ability to reason about preference knowledge represents a fundamental constraint of current RLHF approaches.
From a theoretical perspective, a primary challenge lies in relaxing underlying assumptions. This requires balancing assumptions that are not overly restrictive for practical use cases while maintaining feasibility of theoretical guarantees for computationally efficient algorithms. Key questions include whether algorithms can avoid actively maintaining a policy space and eliminate sub-optimal policies, or relying on computation oracles. Recent work, such as by Wang et al. (2023b) or Wu & Sun (2024), gives hope that this may be possible.
Beyond technical limitations, RLHF raises important ethical and social considerations that require careful attention. These include fundamental questions about whose preferences should guide alignment, ensuring ethical data collection practices that respect privacy, mitigating the amplification of human biases, and addressing incentives for labeler manipulation (Armstrong et al., 2020; Carroll et al., 2023) inherent in current training methodologies. While some technical solutions are emerging, such as privacy-preserving alignment through differential privacy techniques (Wu et al., 2024), addressing these concerns comprehensively requires coordinated progress across technical research, ethical frameworks, and governance structures.
9.3 Conclusion